# Intermittency and non-universality of pair dispersion in isothermal compressible turbulence

### v2

@@ Abstract
Statistical properties of the pair dispersion of Lagrangian particles (tracers) in incompressible turbulent flows provide insights into transport and mixing. We explore the same in transonic to supersonic compressible turbulence of an isothermal ideal gas in two dimensions, driven by large-scale solenoidal and irrotational stirring forces, via direct numerical simulations. We find that the scaling exponents of the order-$p$ negative moments of the distribution of exit times – in particular, the doubling and halving times of pair separations – are nonlinear functions of $p$. Furthermore, the doubling and halving time statistics are different. The halving-time exponents are universal – they satisfy their multifractal model-based prediction, irrespective of the nature of the stirring. However, the doubling-time exponents are not. In the solenoidally-stirred flows, the doubling time exponents can be expressed solely in terms of the multifractal scaling exponents obtained from the structure functions of the solenoidal component of the velocity. Moreover, they depend strongly on the Mach number, Ma, as elongated patches of high vorticity emerge along shock fronts at high Ma. In contrast, in the irrotationally-stirred flows, the doubling-time exponents do not satisfy any known multifractal model-based relation, and are independent of Ma. Our findings are of potential relevance to astrophysical disks and molecular clouds wherein turbulent transport and mixing of gases often govern chemical kinetics and the rates of formation of stars and planetesimals.
keywords:
@@
1 Introduction
[…]
We begin by summarizing the evidence in support of Richardson’s law in incompressible turbulence. Theoretically, it can be understood as a modeling of the evolution of $R(t)$ by a diffusive process (Richardson, 1926; Benzi & Toschi, 2023) with a diffusion constant, $K(R)\sim R^{4/3}$, that stems (Benzi & Toschi, 2023) from Kolmogorov’s 1941 (K41) theory (Kolmogorov, 1941). An alternative interpretation of Richardson’s law is in terms of a dynamic scaling exponent, $z=2/3$, such that typical time scales, $\tau(r)$, of turbulent eddies of size $r$ scale as $\tau(r)\sim r^{z}$. It is now well established, from experiments and DNS, that K41 must be generalized to include intermittency, i.e., the $p$-th moment of the Eulerian velocity difference, $\delta u$, across a length scale $r$ (the order-$p$ velocity structure function), scales as $\left\langle\delta u^{p}(r)\right\rangle\sim r^{\zeta_{p}}$, where $\zeta_{p}$ is a nonlinear concave function of $p$ (Frisch, 1996). This also dictates that turbulence must, in fact, have an infinite number of dynamic exponents (L’vov et al., 1997; Mitra & Pandit, 2004, 2005; Ray et al., 2008, 2011; Pandit et al., 2008), suggesting the presence of intermittency corrections to Richardson’s law, which itself enjoys feeble support from experiments (Ott & Mann, 2000; Shnapp et al., 2018; Elsinga et al., 2022; Tan & Ni, 2022; Shnapp et al., 2023) and direct numerical simulations (DNS) (Salazar & Collins, 2009; Bitane et al., 2012, 2013; Bragg et al., 2016; Buaria et al., 2015; Elsinga et al., 2022) – the observed scaling range of $\left\langle R^{2}\right\rangle$ is at most a decade. A better approach to investigate the intermittency of pair dispersion in turbulence is to use exit times, defined as the time, $\tau_{\alpha}(R)$, it takes for $R$ to cross $\alpha R$ for the first time, where $\alpha$ is fixed. The choice $\alpha>1$ corresponds to the doubling times [^1] To be accurate, $\alpha=2$ means “doubling”. Nevertheless, in practice, any $\alpha>1$ is used (Boffetta & Sokolov, 2002a), $\tau_{\rm D}$, for which (Boffetta et al., 1999; Boffetta & Sokolov, 2002a, b)
@@ 1 Introduction
This bridge relation between the dynamic exponents, $\chi^{\rm D}_{\rm p}$, and $\zeta_{p}$ has been obtained by using the multifractal model of turbulence (Parisi & Frisch, 1985; Frisch, 1996).
@@
In compressible flows, in addition to doubling times, we need to define halving times, $\tau_{\rm H}(R)$, for which $\alpha<1$ and the corresponding scaling exponents are $\chi^{\rm H}_{\rm p}$. As the multifractal model does not distinguish between the doubling and halving times, $\chi^{\rm H}_{\rm p}$ are also expected naively to satisfy the same bridge relation,  (1). Although there have been many studies of the statistics of velocity increments in compressible turbulence (Boldyrev et al., 2002b, a; Schmidt et al., 2008; Konstandin et al., 2012; Wang et al., 2012a; Kritsuk et al., 2007; Wang et al., 2017; Jagannathan & Donzis, 2016), none have investigated the intermittency of tracer pair dispersion. We explore this via simulations of forced compressible turbulence of an isothermal ideal gas, governed by the Navier-Stokes equations. We restrict ourselves to two spatial dimensions in order to reach high resolutions.
Two-dimensional (2D) incompressible turbulence is different from its 3D counterpart because, in addition to the kinetic energy, the enstrophy is conserved in the limit of zero viscosity and forcing. Theory, experiments and DNS indicate that $\left\langle R^{2}(t)\right\rangle$ grows exponentially with time in the direct–enstrophy–cascade regime (Bennett, 1984; Rivera & Ecke, 2005; Jullien, 2003; LaCasce & Ohlmann, 2003). In 2D compressible turbulence, there is small-scale dissipation of kinetic energy via acoustic waves (Falkovich & Kritsuk, 2017). It has been shown that $\left\langle R^{2}(t)\right\rangle$ grows as a power–law beyond the dissipation scales in models of such flows, e.g., surface flows (Cressman & Goldburg, 2003; Li et al., 2024) and the 2D Burgers equation (De et al., 2024).
@@
Here too, we unveil a power-law growth of $\left\langle R^{2}(t)\right\rangle$ across similar spatial scales. While, $\chi^{\rm H}_{\rm p}$ obey the multifractal model-based bridge relation and are universal, $\chi^{\rm D}_{\rm p}\neq\chi^{\rm H}_{\rm p}$, in disagreement with the multifractal model. If the external force is solenoidal: (a) $\chi^{\rm D}_{\rm p}=p-\zeta_{p}^{\rm s}$, where $\zeta_{p}^{\rm s}$ are the scaling exponents of structure functions of the solenoidal component of the velocity field; and (b) the values of $\chi^{\rm D}_{\rm p}$ (and consequently, $\zeta_{p}^{\rm s}$) depend on Ma. This indicates that vortical structures play the dominant role in pair dispersion, and we demonstrate that their fractal dimensions vary significantly with Ma, accounting for the Ma-dependence of $\chi^{\rm D}_{\rm p}$. If the force is purely irrotational, $\chi^{\rm D}_{\rm p}$ neither follows any known bridge relation, nor depends on Ma. In this case, the regions of large flow divergence affect pair dispersion.
In Section 2, we describe the model, and outline its parameters and the details of its DNS. In Section 3, we delineate our analyses of particle pair-dispersion and Lagrangian exit-times, and then provide the concluding remarks in Section 4.
2 Model and its numerical simulation
The equations (Haugen et al., 2022) governing the dynamics of a compressible isothermal ideal gas are:
$\partial_{t}\rho+\partial_{j}(\rho u_{j})$ $=0\,,$ (2a)
$\partial_{t}u_{i}+u_{j}\partial_{j}u_{i}$ $=-\frac{1}{\rho}\partial_{i}P+\frac{1}{\rho}\partial_{j}(\nu\rho\sigma_{ij}+\rho\zeta_{\rm bulk}\delta_{ij}\partial_{k}u_{k})+\mathcal{P}_{ij}(\xi)f_{j}\,,$ (2b)
where $\bm{u}({\bm{x}},t)$ is the velocity field, $\rho({\bm{x}},t)$ is the density field, $P$ is the pressure, $\nu$ is the kinematic shear viscosity, $\zeta_{\rm bulk}$ is the bulk viscosity, $\delta_{ij}$ is the Kronecker delta, $\mathcal{P}_{ij}(\xi)f_{j}$ is the external force, $\sigma_{ij}=(1/2)\left(\partial_{j}u_{i}+\partial_{i}u_{j}\right)-(1/2)\delta_{ij}\partial_{k}u_{k}$ is the traceless rate-of-strain tensor, and the equation of state is given by $P=c_{\rm s}^{2}\rho$, where $c_{\rm s}$ is the (constant) speed of sound. The quantity, ${\bm{f}}$, which we call external force or stirring, is white-in-time and its Fourier transform has compact support at $|{\bm{k}}|=k_{\rm f}=3$, where ${\bm{k}}$ is the wave vector. The Fourier transform of the projection operator, $\mathcal{P}_{ij}(\xi)$ is, $\hat{\mathcal{P}}_{ij}(\xi)\equiv\xi\delta_{ij}+(1-2\xi)k_{i}k_{j}/k^{2}$. We consider two cases, $\xi=0$ and $1$, which correspond, respectively, to an irrotational and a solenoidal external force (see Table 1). The assumption of isothermality holds when the medium is optically thin (transparent to radiation) in which case the heat generated by dissipation at shocks is rapidly radiated away. This is generally valid for molecular clouds which have low density and solar metallicity (Mac Low & Klessen, 2004).
We perform pseudospectral DNS of Eq.  (2) on a 2D doubly-periodic square domain of side $L=2\pi$ with $N^{2}=4096^{2}$ collocation points. The bulk viscosity follows the prescription outlined by Von Neumann & Richtmyer (1950): it is zero everywhere except near the shocks. We assume $\nu$ to be a constant. This is a common practice; see, e.g., Haugen et al. (2022); Federrath et al. (2011); Richard et al. (2022). After the flow reaches a statistically stationary state we seed it uniformly with $n=N^{2}$ Lagrangian particles or tracers at every grid point, such that the distance between a pair of nearest-neighbor tracers is the grid spacing. The position of the $\rm{m}$-th tracer, $\bm{X}^{\rm m}(t)$, evolves according to
$\frac{d\bm{X}^{\rm m}(t)}{dt}={\bm{V}}^{\rm m}(t)\quad\text{where}\quad{\bm{V}}^{\rm m}(t)=\bm{u}({\bm{x}},t)\delta({\bm{x}}-\bm{X}^{\rm m})\/.$ (3)
They cluster at the shocks and their number density maps $\rho({\bm{x}},t)$.

### v3

@@ Abstract
We study statistical properties of the pair dispersion of Lagrangian particles in transonic to supersonic compressible turbulence of an isothermal ideal gas in two dimensions, driven by large-scale solenoidal and irrotational stirring forces, via direct numerical simulations. We find that the scaling exponents of the order-$p$ negative moments of the doubling and halving times of pair separations are nonlinear functions of $p$. Furthermore, the doubling and halving time statistics are different. The halving-time exponents are universal – they satisfy their multifractal model-based prediction, irrespective of the nature of the stirring. The doubling-time exponents are not. In the solenoidally-stirred flows, the doubling time exponents can be expressed solely in terms of the multifractal scaling exponents obtained from the structure functions of the solenoidal component of the velocity. Moreover, they depend strongly on the Mach number, Ma. In contrast, in the irrotationally-stirred flows, the doubling-time exponents do not satisfy any known multifractal model-based relation, and are independent of Ma. Our work is a generalization of Richardson’s law of pair dispersion to compressible turbulence.
@@
One of the pillars of our understanding of turbulence is Richardson’s law [1]: $\left\langle R^{2}(t)\right\rangle\sim t^{3}$, where $R$ is the separation between a pair of Lagrangian particles or tracers, $t$ is time, and $\left\langle\cdot\right\rangle$ denotes averaging over the statistically stationary turbulent state. This law holds for incompressible flows if $R$ lies within the inertial range of length scales, across which the nonlinear interactions dominate over viscous dissipation. Here, we seek a generalization of Richardson’s law to compressible turbulent flows, which are prevalent in many astrophysical systems [2, 3, 4].
We begin by summarizing the present status of Richardson’s law in incompressible turbulence. Theoretically, it can be understood as a modeling of the evolution of $R(t)$ by a diffusive process [1, 5] with a diffusion constant, $K(R)\sim R^{4/3}$, that stems [5] from Kolmogorov’s 1941 (K41) theory [6]. Stated differently, Richardson’s law implies that there exists a dynamic scaling exponent, $z=2/3$, such that typical time scales, $\tau(r)$, of turbulent eddies of size $r$ scale as $\tau(r)\sim r^{z}$. It is now well established, from experiments and direct numerical simulations (DNS), that K41 must be generalized to include intermittency, i.e., the $p$-th moment of the Eulerian velocity difference, $\delta u$, across a length scale $r$ (the order-$p$ velocity structure function), scales as $\left\langle\delta u^{p}(r)\right\rangle\sim r^{\zeta_{p}}$, where $\zeta_{p}$ is a nonlinear concave function of $p$ [7]. This also dictates that turbulence must, in fact, have an infinite number of dynamic exponents [8, 9, 10, 11, 12, 13], suggesting the presence of intermittency corrections to Richardson’s law, which itself enjoys feeble support from experiments [14, 15, 16, 17, 18] and DNS [19, 20, 21, 22, 23, 16] – the observed scaling range of $\left\langle R^{2}\right\rangle$ is at most a decade. A better approach to investigate the intermittency of pair dispersion in turbulence is to use exit times, defined as the time, $\tau_{\sigma}(R)$, it takes for $R$ to cross $\sigma R$ for the first time, where $\sigma$ is fixed. The choice $\sigma>1$ corresponds to the doubling times [^1] To be accurate, $\sigma=2$ means “doubling”. Nevertheless, in practice, any $\sigma>1$ is used [26], $\tau_{\rm D}$, for which [25, 26, 27]
@@
This bridge relation between the dynamic exponents, $\chi^{\rm D}_{\rm p}$, and $\zeta_{p}$ has been obtained by using the multifractal model of turbulence.
@@
In compressible flows, in addition to doubling times, we need to define halving times $\tau_{\rm H}(R)$, for which $\sigma<1$ and the corresponding scaling exponents are $\chi^{\rm H}_{\rm p}$. As the multifractal model does not distinguish between the doubling and halving times, $\chi^{\rm H}_{\rm p}$ are also expected naively to satisfy the same bridge relation,  (1). Although there have been many studies of the statistics of velocity increments in compressible turbulence [28, 29, 30, 31, 32, 33] and tracer dispersion in compressible surface flows are known not to obey Richardson’s law [34], the intermittency of tracer pair dispersion have not been studied. We explore this via simulations of forced compressible turbulence of an isothermal ideal gas, governed by the Navier-Stokes equations in two dimensions.
In 2D compressible turbulence forced at an intermediate scale, Refs. [35, 36] had obtained two cascades of the kinetic energy: an inverse cascade from the energy injection scale to large length scales and a forward cascade to small length scales. In what follows, we choose the energy-injection scales to be close to the linear size of our simulation domain; and we focus on the forward- cascade regime, which we show to be intermittent.
Figure 1: Typical snapshots of $\nabla\cdot{\bf u}$, the vorticity ${\bm{\omega}}={\bm{\nabla}}\times\bm{u}$, and $\log_{10}\rho$, for solenoidally-driven supersonic turbulence, run S2 (${\bm{\nabla}}\cdot\bm{u}$ and $\bm{\omega}$} are normalized by their respective root-mean-square values). Thin filaments in the ${\bm{\nabla}}\cdot\bm{u}$ profile are the shocks. The vorticity, $\bm{\omega}$}, is large in the vicinity of curved shocks. There are strong density gradients across shocks.
@@
We unveil a power-law growth of $\left\langle R^{2}(t)\right\rangle$, but the scaling regime is too small to make robust quantitative measurements. Thus, we resort to the exit-time approach to calculate the dynamic exponents $\chi^{\rm H}_{\rm p}$ and $\chi^{\rm D}_{\rm p}$. We find that $\chi^{\rm H}_{\rm p}$ obey the multifractal model-based bridge relation and are universal, i.e., does not depend on how the turbulence in forced. But contrary to the prediction of the multifractal model $\chi^{\rm D}_{\rm p}\neq\chi^{\rm H}_{\rm p}$. Moreoever, $\chi^{\rm D}_{\rm p}$ is not universal. If the external force is solenoidal, then: (a) $\chi^{\rm D}_{\rm p}=p-\zeta_{p}^{\rm s}$, where $\zeta_{p}^{\rm s}$ are the scaling exponents of structure functions of the solenoidal component of the velocity field; and (b) $\chi^{\rm D}_{\rm p}$ depend on the Mach number (Ma). If the force is purely irrotational, $\chi^{\rm D}_{\rm p}$ neither follows any known bridge relation, nor depends on Ma. We further identify, through the multifractal model, the dimensions of the structures that play the dominant role in pair dispersion.
We perform pseudospectral DNS of the equations of compressible, isothermal gas with a random white-in-time external force (at the Fourier modes with $\mid{\bm{k}}\mid=k_{\rm f}=3$) on a 2D doubly-periodic square domain of side $L=2\pi$ with $N^{2}=4096^{2}$ collocation points. The assumption of isothermality holds when the gas is optically thin (transparent to radiation) in which case the heat generated by the dissipation at shocks is rapidly radiated away. This is generally valid for molecular clouds which have low density and solar metallicity [2]. In addition to shear viscosity, we use a bulk viscosity which is zero everywhere except near the shocks [37]. Shocks are generally resolved across $4$ to $10$ grid points. After the flow reaches a statistically stationary state we seed it uniformly with $n=N^{2}$ Lagrangian particles and track them as the flow evolves. They cluster at the shocks and their number density is proportional to $\rho({\bm{x}},t)$.